This is the question that I was given:
Find a number T such that the line passing through the two given point has a slope of -2.
1) (0,t),(9,4)
2) (1,t),(2,4)
3) (t,t),(5,9)

Answer by stanbon(75887)

This is the question that I was given:
Find a number T such that the line passing through the two given point has a slope of -2.
1) (0,t),(9,4)
(t-4)/(0-9) = -2
t-4 = 18
t = 22
------------------------
2) (1,t),(2,4)
(t-4)/(1-2) = -2
t-4 = 2
t = 6
-----------------------
3) (t,t),(5,9)
(t-9)/(t-5) = -2
t-9 = -2(t-5)
t-9 = -2t+10
3t = 19
t = 19/3
==================
Cheers,
Stan H.
====================